OpenStudy (anonymous):

Stephanie is organizing the seating arrangements for a concert in her local park. The first rows seats are modeled by the function f(x)=15. To give everyone a good view, each row much have 4 more seats than the row before it. The number of additional seats a row has is modeled by the function s(x)=4(x-1). Explain to Stephanie how to create an equation to predict the number of seats in any row. Describe how to determine the number of seats in the 40th row. Use complete sentences.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ s(x)+f(x)\huge\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{(s+f)(x)=15+4(x-1)\huge\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{(s+f)(x)=15+4x-4\huge\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{(s+f)(x)=11+4x\huge\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

This is for the number of seats for any Xth row.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{(s+f)(40)=11+4(40)=11+160=171\huge\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

good ?

OpenStudy (anonymous):

Thank you so much @SolomonZelman ! You're a lifsaver!

OpenStudy (solomonzelman):

Anytime!